Summary - 080406

Overall a smooth and productive night. After an initial rebooting around sunset, LBC didn't fail the entire night. There were several small telescope slitches, but the loss of time was not significant. We struggled with focus somewhat, and there were occasionally guiding issues, and clouds were pretty bad for about an hour. Otherwise the night was very smooth. A number of OBs on UMaII, EGS and Bootes. Seeing OK but not excellent, fluctuating between 0.7 and 1.5".

Details

Observers: Xiaohui Fan, Ben Weiner

Supporting Astronomer: Marco Pedani

Operator Aaron Ceranski

All Times UT

0045 has red rotator error in the beginning of bias sequence; restart camera

0049 25Bias_Bino

0110 blue controller error: too few downloaded pixels. 19 biases in blue; 25 red completed. restarting camera

0112 25Bias_Bino again

0113 blue pressure reading zero. stopped again. restart

0133 after another few communication errors, we are rebooting CMU. Meanwhile, light cirrus on the west horizon, so non-photometric tonight.

0145 we are open

0202 SkyFlat_UsY_5; 20000 ADU in Us; 12000 ADU in Y

0212 SkyFlat_UY_5; 19000 in U; 9000 in Y

0221 SkyFlat_Br_5: 33000 in B, 18000 in r

0310 finally converging on focus:

Blue: -268 -150 296 18 307 138 1.44
Red: -32 -139 4 81 62 -62 0.95

0315 UMaII_off8_Br 300sec x 6; 1.5" in B; 1.3" in r

0353 lbcfpia

Blue: -354 168 -262 8 -73 101 1.32 Red: -165 -30 -218 -62 7 -27 0.84

0403 UMaII_off9_Br 300sec x 6; 1.5" in B, 1.3" in r

0440 lbcfpia; seeing seems worse

Blue: 5 -103 337 155 117 0 1.75 Red: -160 215 -168 17 -192 46 1.38

0447 UMaII_off10_Br 300sec x6; 2.0" in B, 1.7" in r; improved to 1.7 in B and 1.4 in r

0527 moving to Bootes; lbcfpia

0553 struggled with focus and variable seeing. finally starting b51r, 240sx5

blue guiding failing, red guiding on faint stars. images look enlongated. 1.5" in Y., and some cirrus as well. stopped after second expo. move to 53r for better luck.

0616 starting b53r 240sx5 Y: 1.2", U_s: 1.7", clouds passing through.

0648 b53r again after lbcfpia 240x5 ; Y: 1.0", U_s: 1.5", still clouds coming and going; in some cases no guide star and images trailed.

0724 b14r 240x5; seeing improving and sky seems clear again. Y: 0.9", U_s: 1.3"; mirror jumped at travel limit during 2rd exp. image #4 trailed as well. made an additional OB b14r_2 to add two random offsets so complete this pointing.

0810 moving to EGS. EGS_Y_3a, 240x5; Y: 0.8", U: 1.3" ; continues to improve.

0845 lbcfpia then EGS_Y_3b, 240x5; Y: 0.8"; U: 1.3"

0926 back to Bootes; b51r 240x5; Y: 0.7", U_s: 1.1"

1007 b11r 240x5, Y:0.8", U_s: 1.3"; last image has large (2000 or so) focus drift

1047 b13r 240x5, Y: 0.8", U_s:1.3", tracking stopping in the middle of the fourth exposure. lbcr.20080406.110321.fits junk. adding an additional expo with random offset (OB: b13r_1)

1130 b47r 240x5: Y: 0.9" increased to 1.3"; add another 240 expo (position 1) during twilight

1216 SkyFlat_Br_5

1226 SkyFlat_UY_rot180_5

1234 SkyFlat_UsI_rot180_5 (I filter saturated)

-- XiaohuiFan - 06 Apr 2008
Topic revision: r5 - 06 Apr 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback